# https://rlwg-grndctrl.github.io/

### v1

@@ Abstract
Recent advances in video world modeling have enabled large-scale generative models to simulate embodied environments with high visual fidelity, providing strong priors for prediction, planning, and control. Yet, despite their realism, these models often lack geometric grounding, limiting their use in navigation tasks that require spatial coherence and long-horizon stability. We introduce Reinforcement Learning with World Grounding (RLWG), a self-supervised post-training framework that aligns pretrained world models with a physically verifiable structure through geometric and perceptual rewards. Analogous to reinforcement learning from verifiable feedback (RLVR) in language models, RLWG can use multiple rewards that measure pose cycle-consistency, depth reprojection, and temporal coherence. We instantiate this framework with GrndCtrl, a reward-aligned adaptation method based on Group Relative Policy Optimization (GRPO), yielding world models that maintain stable trajectories, consistent geometry, and reliable rollouts for embodied navigation. Like post-training alignment in large language models, GrndCtrl leverages verifiable rewards to bridge generative pretraining and grounded behavior, achieving superior spatial coherence and navigation stability over supervised fine-tuning in outdoor environments.
@@ 1 Introduction
Large-scale video world models have emerged as powerful priors for modeling perception and control for embodied agents [nvidia2025cosmosworldfoundationmodel, bar2025navigationworldmodels, assran2025vjepa2, jin2025posepilotsteeringcamerapose, ren2025gen3c]. By learning to predict future observations from past frames and actions, these models approximate the transition dynamics of the physical world, enabling simulation, planning, and policy evaluation. Operating in the pixel domain aligns them with real-world sensors and exploits the vast implicit supervision available in video, allowing unified modeling across domains such as manipulation, driving, and navigation. Yet despite their impressive generative fidelity, these models are often incentivized to capture the appearance of motion more than its structure. Their rollouts remain visually plausible but geometrically and temporally inconsistent: poses drift, depths wobble, and trajectories lose alignment over time. Even subtle deviations in inferred geometry accumulate into compounding spatial errors corrupting metric structure. These instabilities limit the use of current models for closed-loop tasks such as localization, mapping, and planning, where physically consistent representation is essential.
@@
We define world model grounding as aligning learned dynamics with physically verifiable spatial and temporal invariants, so that rollouts honor geometry and time in addition to reproducing surface appearance. Grounding shifts the objective of world modeling from visual plausibility to structural consistency, ensuring that the model’s internal dynamics respect the constraints of real motion and scene structure. To this end, we introduce Reinforcement Learning with World Grounding (RLWG), a self-supervised post-training framework that refines pretrained world models using verifiable geometric and perceptual rewards derived from model rollouts. RLWG extends the principle of Reinforcement Learning with Verifiable Rewards (RLVR) from language models [lambert2025tulu3pushingfrontiers] to the embodied domain, replacing text-based logical verification with geometric and temporal verification. In RLWG, a pretrained world model is treated as a policy that generates multiple candidate rollouts from the same context; each rollout is automatically scored using verifiable grounding rewards that quantify spatial and temporal coherence, such as pose cycle-consistency, depth reprojection agreement and action adherence. Unlike reconstruction losses that only penalize pixel error, these rewards measure physical correctness of the rollouts.
@@
To optimize these verifiable rewards efficiently, we adopt Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] as our training mechanism, yielding our algorithm, GrndCtrl. For each context (and actions when available), the model generates a group of rollouts that are ranked by their grounding rewards; relative advantages are computed within the group, and the latent transition operator is updated using a clipped policy gradient objective regularized toward the pretrained model. This formulation preserves visual quality while progressively aligning the model’s dynamics with measurable structure in the real world. The process requires no human annotations or external simulators, operating entirely through self-supervised reinforcement grounded in the model’s own predictions. Conceptually, GrndCtrl extends the success of GRPO-based alignment in generative modeling to the geometric domain, grounding visual world models in verifiable 3D and temporal coherence.
@@ 2.1 Controlling World Models
Recent progress in large-scale video foundation models has transformed video prediction into controllable world simulation. Models such as Cosmos-Predict [nvidia2025cosmosworldfoundationmodel], and V-JEPA [assran2025vjepa2] unify multi-modal conditioning for long-horizon prediction and control. These models achieve impressive simulation fidelity but still exhibit spatial drift, geometric misalignment, and temporal incoherence over extended rollouts, revealing limitations in geometric grounding. Architectural innovations like flow matching, conditional diffusion transformers, and masked latent prediction have improved realism but not physical consistency.
Controlling these models can be categorized into action-conditioned and camera-conditioned paradigms. The first paradigm trains models to predict futures from discrete actions, and can be further defined by the nature of the action and observation frame. Some methods [zhou2025dinowmworldmodelspretrained, zhu2025unifiedworldmodelscoupling] assume a static camera observing an embodiment performing actions, while others [hafner2024masteringdiversedomainsworld, valevski2025diffusionmodelsrealtimegame, bar2025navigationworldmodels, bai2025wholebodyconditionedegocentricvideo] assume a fixed camera on the embodiment and train models to predict ego-centric views. Jointly, they aim to enable model-predictive control, but face challenges with physical realism. The second paradigm explicitly decouples viewpoint from embodiment, allowing the model to generate future observations from arbitrary camera poses. Early works [wang2024motionctrlunifiedflexiblemotion, he2025cameractrlenablingcameracontrol] injected pose embeddings into diffusion models to achieve this control, but lacked explicit geometric alignment. Subsequent methods [jin2025posepilotsteeringcamerapose, ren2025gen3c] improved consistency via self-supervised warping, 3D-informed point cloud conditioning, achieving more precise viewpoint control. Recent work [guo2025ctrlworldcontrollablegenerativeworld] bridges the two paradigms by jointly training on multiple viewpoints, and improved spatial-temporal consistency with pose-conditioned memory retrieval. Despite these advances, most approaches rely on supervised learning or one-step consistency objectives and remain open-loop, with no mechanism to evaluate or optimize physical correctness.
@@ 2.2 Reward Learning for Post-Training
Reinforcement-based post-training has become central to aligning large generative models. In language systems, Reinforcement Learning from Human Feedback (RLHF) [ouyang2022traininglanguagemodelsfollow] and Reinforcement Learning with Verifiable Rewards (RLVR) [lambert2025tulu3pushingfrontiers] replace imitation with objective-driven alignment, while Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] stabilizes learning via stochastic rollouts comparative updates. Extensions to vision, such as DanceGRPO [xue2025dancegrpounleashinggrpovisual], demonstrate that rewards on visual quality can fine-tune video diffusion models effectively.
Building on this foundation, RLWG adapts RLVR to world modeling, optimizing pretrained world models using physically verifiable rewards including cycle-consistency, depth reprojection, and trajectory stability. GrndCtrl instantiates RLWG as a multi-objective GRPO over grounded rewards. This process enforces geometric coherence without human supervision, significantly reducing long-horizon drift. Parallel advances in 3D perception emphasize similar constraints: VGGT [wang2025vggt] and MapAnything [keetha2025mapanythinguniversalfeedforwardmetric] predict depth and camera pose for consistent scene reconstruction, while SpaTracker [xiao2025spatialtrackerv23dpointtracking] integrates rigidity priors for robust 3D tracking. Together, these efforts point toward reward-informed geometry as a unifying principle, where physical correctness acts as an alignment signal bridging generative modeling, simulation, and control.
@@ 3.2 Verifiable Self-Supervised Rewards
We adopt the depth inlier ratio from MapAnything [keetha2025mapanythinguniversalfeedforwardmetric] evaluations as a verifiable geometric reward for depth temporal reprojection. For each pixel $p\!\in\!\Omega$, define the reprojected correspondence and expected depth via the evaluator geometry
@@
where $\Phi$ back-projects $p$ using $D_{t}$, applies $(\Delta R_{t},\Delta\mathbf{t}_{t})$, and projects into frame t+1t{+}1. The per-pair depth inlier ratio (threshold $\gamma{=}0.0103$) is
@@
We use the frozen VideoAlign [liu2025improving] as evaluator $\mathcal{V}$, which returns three sequence-level scores in [0,1][0,1]—visual quality $s_{\text{vis}}$, motion quality $s_{\text{mot}}$, and text alignment $s_{\text{txt}}$—for a rollout $\hat{x}_{1:T}$ (optionally conditioned on a prompt $y$):
@@ 3.3 GRPO for RLWG Post-Training
Groups are instantiated by sampling multiple candidate rollouts from the diffusion generator under controlled stochasticity. Diffusion sampling can be formulated as a reverse-time stochastic differential equation (SDE) or, equivalently in time-marginals, as a probability–flow ordinary differential equation (ODE) [song2021scorebasedgenerativemodelingstochastic, albergo2025stochasticinterpolantsunifyingframework]. Let $x_{t}\in\mathbb{R}^{d}$ denote the latent state at time $t\in[0,1]$ (noisiest at t=1t{=}1, data manifold at $t{\to}0$). A standard variance–preserving forward SDE is
@@ 4.1 Datasets
We train and evaluate our methods on three datasets spanning diverse embodiments and scenarios. CODa [zhang2023towards] is a campus navigation dataset collected on wheeled robots with pseudo-ground truth poses. SCAND [karnan2022sociallycompliantnavigationdataset] is a social navigation dataset also collected on campus, featuring embodiments of both a wheeled robot and a quadruped. CityWalk [liu2025citywalker] is an egocentric urban navigation dataset of a person walking in crowded city streets, collected from in-the-wild YouTube city walking videos and reprocessed with MapAnything [keetha2025mapanythinguniversalfeedforwardmetric] to obtain pose estimates. We subsample each dataset to 2k non-overlapping 13-frame sequences of pose-action pairs for training, ensuring temporal diversity and avoiding data leakage.
@@ 4.2 Baseline Model
We obtain the baseline pretrained world model $W_{\theta}$ via supervised fine-tuning (SFT) on Cosmos-Predict2-2B-Video2World [nvidia2025cosmosworldfoundationmodel], a diffusion model with a latent VAE backbone and temporal attention. We use a modified version of its action-conditioned video predictor post-training pipeline, adapting the action space from 7 to 6 degrees of freedom to match our navigation setting: $(x,y,z,\mathrm{roll},\mathrm{pitch},\mathrm{yaw})$ as 6×12 action embeddings. We initialize the action embeddings randomly while fine-tuning the entire DiT backbone for 20k steps with an effective batch size of 64. Training follows the EDM framework [karras2022elucidatingdesignspacediffusionbased, karras2024analyzingimprovingtrainingdynamics], using the weighted expectation of denoising score matching loss over noise levels. We perform full SFT over all DiT backbone parameters keeping the visual encoder and decoder frozen, as our experiments showed that only full SFT or full-size LoRA yielded meaningful changes in motion dynamics. Additional implementation details are provided in Appendix.
@@ 4.3 GRPO Post-Training
Post-training applies GRPO with self-supervised verifiable rewards as described in Sec. 3.3. We use MapAnything [keetha2025mapanythinguniversalfeedforwardmetric] as $\mathcal{E}$ to obtain rewards $_{\text{trans}}, rr$$_{\text{rot}}$, and $r_{\text{dtr}}$. We use VideoAlign [liu2025improving] as $\mathcal{V}$ to obtain the reward $r_{\text{v}}$. We perform an ablation study of combinatorial rewards as defined in Sec. 3.2 to evaluate multi-objective GRPO. For each reward configuration, we train for 100 steps with an effective batch size of 8, generating $G=8$ stochastic rollouts $\{\hat{x}^{(i)}_{1:T}\}_{i=1}^{G}$ per conditioning context $c=(x_{0},a_{0:T-1})$. To obtain diverse rollouts, we use the same initial noise for reverse-SDE diffusion with the same starting frame $x_{0}$ and action trajectory $a_{0:T-1}$, but inject stochastic Brownian noise at each diffusion step as described in Sec. 3.3. We compute per-step likelihood ratios only over the first 60% of diffusion timesteps to focus training on the most relevant denoising steps to improve training stability. Additional training details are provided in Appendix.
@@ 5 Analysis and Discussion
CODa [zhang2023towards]
Baseline [nvidia2025cosmosworldfoundationmodel] $57.8$ $1.77$ $7.40$ $38.9$ $71.5$ $1.55$ $7.41$ $39.1$ $56.9$ $1.71$ $7.40$ $38.3$
@@
SCAND [karnan2022sociallycompliantnavigationdataset]
Baseline [nvidia2025cosmosworldfoundationmodel] $186.3$ $3.76$ $7.16$ $23.6$ $315.9$ $4.24$ $7.13$ $21.4$ $117.0$ $4.02$ $6.99$ $18.4$
@@
CityWalk [liu2025citywalker]
Baseline [nvidia2025cosmosworldfoundationmodel] $11.7$ $3.13$ $7.96$ $46.9$ $13.1$ $3.27$ $7.94$ $47.4$ $20.8$ $4.47$ $7.90$ $44.5$
@@ 5.1 Impact of GrndCtrl on World Model Failures
Incorporating perceptual feedback through the full reward set (T+R+DTRI+V) produces the most balanced and robust performance across all datasets. On CODa, the full objective achieves translation error reductions of $31\%$ (Seen), $43\%$ (Counterfactual), and $45\%$ (Unseen) relative to baseline, with rotation error improving by $28\%$ (Seen), $8\%$ (Counterfactual), and $11\%$ (Unseen). On SCAND, translation error reduces by $28\%$ (Seen), $30\%$ (Counterfactual), and $5\%$ (Unseen), while CityWalk shows consistent improvements with Counterfactual achieving a $64\%$ reduction. The full objective recovers and extends the rollout accuracy of T+R while preserving local stability from DTRI. Perceptual alignment via video-based evaluators acts as a long-horizon stabilizer, promoting rollouts that are both physically consistent and visually coherent. Figure 3 showcases qualitative comparisons of GrndCtrl rollouts $\hat{x}_{1:T}$ against the baseline. GrndCtrl significantly mitigates the scene drift failures frequently observed when queried with mirrored input action sequences $a_{0:T-1}$, and improves rollout consistency.

### v2

@@ Abstract
Recent advances in video world modeling have enabled large-scale generative models to simulate embodied environments with high visual fidelity, providing strong priors for prediction, planning, and control. Yet, despite their realism, these models often lack geometric grounding, limiting their use in navigation tasks that require spatial coherence and stability. We introduce Reinforcement Learning with World Grounding (RLWG), a self-supervised post-training framework that aligns pretrained world models with a physically verifiable structure through geometric and perceptual rewards. Analogous to reinforcement learning from verifiable feedback (RLVR) in language models, RLWG can use multiple rewards that measure pose cycle-consistency, depth reprojection, and temporal coherence. We instantiate this framework with GrndCtrl, a reward-aligned adaptation method based on Group Relative Policy Optimization (GRPO), yielding world models that maintain stable trajectories, consistent geometry, and reliable rollouts for embodied navigation. Like post-training alignment in large language models, GrndCtrl leverages verifiable rewards to bridge generative pretraining and grounded behavior, achieving superior spatial coherence and navigation stability over supervised fine-tuning in outdoor environments.
@@
00footnotetext: Work done while intern at FieldAI.
@@ 1 Introduction
Large-scale video world models have emerged as powerful priors for modeling perception and control for embodied agents [17, 4, 2, 9, 19]. By learning to predict future observations from past frames and actions, these models approximate the transition dynamics of the physical world, enabling simulation, planning, and policy evaluation. Operating in the pixel domain aligns them with real-world sensors and exploits the vast implicit supervision available in video, allowing unified modeling across domains such as manipulation, driving, and navigation. Yet despite their impressive generative fidelity, these models are often incentivized to capture the appearance of motion more than its structure. Their rollouts remain visually plausible but geometrically and temporally inconsistent: poses drift, depths wobble, and trajectories lose alignment over time. Even subtle deviations in inferred geometry accumulate into compounding spatial errors corrupting metric structure. These instabilities limit the use of current models for closed-loop tasks such as localization, mapping, and planning, where physically consistent representation is essential.
@@
We define world model grounding as aligning learned dynamics with physically verifiable spatial and temporal invariants, so that rollouts honor geometry and time in addition to reproducing surface appearance. Grounding shifts the objective of world modeling from visual plausibility to structural consistency, ensuring that the model’s internal dynamics respect the constraints of real motion and scene structure. To this end, we introduce Reinforcement Learning with World Grounding (RLWG), a self-supervised post-training framework that refines pretrained world models using verifiable geometric and perceptual rewards derived from model rollouts. RLWG extends the principle of Reinforcement Learning with Verifiable Rewards (RLVR) from language models [14] to the embodied domain, replacing text-based logical verification with geometric and temporal verification. In RLWG, a pretrained world model is treated as a policy that generates multiple candidate rollouts from the same context; each rollout is automatically scored using verifiable grounding rewards that quantify spatial and temporal coherence, such as pose cycle-consistency, depth reprojection agreement and action adherence. Unlike reconstruction losses that only penalize pixel error, these rewards measure physical correctness of the rollouts.
@@
To optimize these verifiable rewards efficiently, we adopt Group Relative Policy Optimization (GRPO) [21] as our training mechanism, yielding our algorithm, GrndCtrl. For each context (and actions when available), the model generates a group of rollouts that are ranked by their grounding rewards; relative advantages are computed within the group, and the latent transition operator is updated using a clipped policy gradient objective regularized toward the pretrained model. This formulation preserves visual quality while progressively aligning the model’s dynamics with measurable structure in the real world. The process requires no human annotations or external simulators, operating entirely through self-supervised reinforcement grounded in the model’s own predictions. Conceptually, GrndCtrl extends the success of GRPO-based alignment in generative modeling to the geometric domain, grounding visual world models in verifiable 3D and temporal coherence.
@@ 2.1 Controlling World Models
Recent progress in large-scale video foundation models has transformed video prediction into controllable world simulation. Models such as Cosmos-Predict [17], and V-JEPA [2] unify multi-modal conditioning for long-horizon prediction and control. These models achieve impressive simulation fidelity but still exhibit spatial drift, geometric misalignment, and temporal incoherence over extended rollouts, revealing limitations in geometric grounding. Architectural innovations like flow matching, conditional diffusion transformers, and masked latent prediction have improved realism but not physical consistency.
Controlling these models can be categorized into action-conditioned and camera-conditioned paradigms. The first paradigm trains models to predict futures from discrete actions, and can be further defined by the nature of the action and observation frame. Some methods [29, 30] assume a static camera observing an embodiment performing actions, while others [6, 23, 4, 3] assume a fixed camera on the embodiment and train models to predict ego-centric views. Jointly, they aim to enable model-predictive control, but face challenges with physical realism. The second paradigm explicitly decouples viewpoint from embodiment, allowing the model to generate future observations from arbitrary camera poses. Early works [25, 7] injected pose embeddings into diffusion models to achieve this control, but lacked explicit geometric alignment. Subsequent methods [9, 19] improved consistency via self-supervised warping, 3D-informed point cloud conditioning, achieving more precise viewpoint control. Recent work [5] bridges the two paradigms by jointly training on multiple viewpoints, and improved spatial-temporal consistency with pose-conditioned memory retrieval. Despite these advances, most approaches rely on supervised learning or one-step consistency objectives and remain open-loop, with no mechanism to evaluate or optimize physical correctness.
@@ 2.2 Reward Learning for Post-Training
Reinforcement-based post-training has become central to aligning large generative models. In language systems, Reinforcement Learning from Human Feedback (RLHF) [18] and Reinforcement Learning with Verifiable Rewards (RLVR) [14] replace imitation with objective-driven alignment, while Group Relative Policy Optimization (GRPO) [21] stabilizes learning via stochastic rollouts comparative updates. Extensions to vision, such as DanceGRPO [27], demonstrate that rewards on visual quality can fine-tune video diffusion models effectively.
Building on this foundation, RLWG adapts RLVR to world modeling, optimizing pretrained world models using physically verifiable rewards including cycle-consistency, depth reprojection, and trajectory stability. GrndCtrl instantiates RLWG as a multi-objective GRPO over grounded rewards. This process enforces geometric coherence without human supervision, significantly reducing scene drift. Parallel advances in 3D perception emphasize similar constraints: VGGT [24] and MapAnything [13] predict depth and camera pose for consistent scene reconstruction, while SpaTracker [26] integrates rigidity priors for robust 3D tracking. Together, these efforts point toward reward-informed geometry as a unifying principle, where physical correctness acts as an alignment signal bridging generative modeling, simulation, and control.
@@ 3.2 Verifiable Self-Supervised Rewards
We adopt the depth inlier ratio from MapAnything [13] evaluations as a verifiable geometric reward for depth temporal reprojection. For each pixel $p\!\in\!\Omega$, define the reprojected correspondence and expected depth via the evaluator geometry
@@
where $K_{t},K_{t+1}$ are the intrinsics of frames $t$ and t+1t{+}1, and $d_{t\to t+1}^{\exp}(p)$ is the depth at t+1t{+}1 obtained by back-projecting $p$ with $D_{t}(p)$, transforming by $(\Delta R_{t},\Delta\mathbf{t}_{t})$, and re-projecting to $\hat{p}$. The per-pair depth inlier ratio (threshold $\gamma{=}0.0103$) is
@@
We use the frozen VideoAlign [15] as evaluator $\mathcal{V}$, which returns three sequence-level scores in [0,1][0,1]—visual quality $s_{\text{vis}}$, motion quality $s_{\text{mot}}$, and text alignment $s_{\text{txt}}$—for a rollout $\hat{x}_{1:T}$ (optionally conditioned on a prompt $y$):
@@ 3.3 GRPO for RLWG Post-Training
Groups are instantiated by sampling multiple candidate rollouts from the diffusion generator under controlled stochasticity. Diffusion sampling can be formulated as a reverse-time stochastic differential equation (SDE) or, equivalently in time-marginals, as a probability–flow ordinary differential equation (ODE) [22, 1]. Let $x_{t}\in\mathbb{R}^{d}$ denote the latent state at time $t\in[0,1]$ (noisiest at t=1t{=}1, data manifold at $t{\to}0$). A standard variance–preserving forward SDE is
@@ 4.1 Datasets
We train and evaluate our methods on three datasets spanning diverse embodiments and scenarios. CODa [28] is a campus navigation dataset collected on wheeled robots with pseudo-ground truth poses. SCAND [10] is a social navigation dataset also collected on campus, featuring embodiments of both a wheeled robot and a quadruped. CityWalk [16] is an egocentric urban navigation dataset of a person walking in crowded city streets, collected from in-the-wild YouTube city walking videos and reprocessed with MapAnything [13] to obtain pose estimates. We subsample each dataset to 2k non-overlapping 13-frame sequences of pose-action pairs for training, ensuring temporal diversity and avoiding data leakage.
@@ 4.2 Baseline Model
We obtain the baseline pretrained world model $W_{\theta}$ via supervised fine-tuning (SFT) on Cosmos-Predict2-2B-Video2World [17], a diffusion model with a latent VAE backbone and temporal attention. We use a modified version of its action-conditioned video predictor post-training pipeline, adapting the action space from 7 to 6 degrees of freedom to match our navigation setting: $(x,y,z,\mathrm{roll},\mathrm{pitch},\mathrm{yaw})$ as 6×12 action embeddings. We initialize the action embeddings randomly while fine-tuning the entire DiT backbone for 20k steps with an effective batch size of 64. Training follows the EDM framework [11, 12], using the weighted expectation of denoising score matching loss over noise levels. We perform full SFT over all DiT backbone parameters keeping the visual encoder and decoder frozen, as our experiments showed that only full SFT or full-size LoRA yielded meaningful changes in motion dynamics. Additional implementation details are provided in Appendix.
@@ 4.3 GRPO Post-Training
Post-training applies GRPO with self-supervised verifiable rewards as described in Sec. 3.3. We use MapAnything [13] as $\mathcal{E}$ to obtain rewards $_{\text{trans}}, rr$$_{\text{rot}}$, and $r_{\text{dtr}}$. We use VideoAlign [15] as $\mathcal{V}$ to obtain the reward $r_{\text{v}}$. We perform an ablation study of combinatorial rewards as defined in Sec. 3.2 to evaluate multi-objective GRPO. For each reward configuration, we train for 100 steps with an effective batch size of 8, generating $G=8$ stochastic rollouts $\{\hat{x}^{(i)}_{1:T}\}_{i=1}^{G}$ per conditioning context $c=(x_{0},a_{0:T-1})$. To obtain diverse rollouts, we use the same initial noise for reverse-SDE diffusion with the same starting frame $x_{0}$ and action trajectory $a_{0:T-1}$, but inject stochastic Brownian noise at each diffusion step as described in Sec. 3.3. We compute per-step likelihood ratios only over the first 60% of diffusion timesteps to focus training on the most relevant denoising steps to improve training stability. Additional training details are provided in Appendix.
@@ 5 Analysis and Discussion
CODa [28]
Baseline [17] $57.8$ $1.77$ $7.40$ $38.9$ $71.5$ $1.55$ $7.41$ $39.1$ $56.9$ $1.71$ $7.40$ $38.3$
@@
SCAND [10]
Baseline [17] $186.3$ $3.76$ $7.16$ $23.6$ $315.9$ $4.24$ $7.13$ $21.4$ $117.0$ $4.02$ $6.99$ $18.4$
@@
CityWalk [16]
Baseline [17] $11.7$ $3.13$ $7.96$ $46.9$ $13.1$ $3.27$ $7.94$ $47.4$ $20.8$ $4.47$ $7.90$ $44.5$
@@ 5.1 Impact of GrndCtrl on World Model Failures
Incorporating perceptual feedback through the full reward set (T+R+DTRI+V) produces the most balanced and robust performance across all datasets. On CODa, the full objective achieves translation error reductions of $31\%$ (Seen), $43\%$ (Counterfactual), and $45\%$ (Unseen) relative to baseline, with rotation error improving by $28\%$ (Seen), $8\%$ (Counterfactual), and $11\%$ (Unseen). On SCAND, translation error reduces by $28\%$ (Seen), $30\%$ (Counterfactual), and $5\%$ (Unseen), while CityWalk shows consistent improvements with Counterfactual achieving a $64\%$ reduction. The full objective recovers and extends the rollout accuracy of T+R while preserving local stability from DTRI. Perceptual alignment via video-based evaluators acts as a stabilizer, promoting rollouts that are both physically consistent and visually coherent. Figure 3 showcases qualitative comparisons of GrndCtrl rollouts $\hat{x}_{1:T}$ against the baseline. GrndCtrl significantly mitigates the scene drift failures frequently observed when queried with mirrored input action sequences $a_{0:T-1}$, and improves rollout consistency.